OpenStudy (anonymous):

HEEEEEEELLLPPP!!!!! The lengths of the sides of a rectangle become 3 times greater than the original measurements. Which statement is true about the rectangle?

OpenStudy (compassionate):

Give us the statements...

OpenStudy (anonymous):

Its area becomes 9 times greater, and perimeter becomes 3 times greater. Its area becomes 3 times greater, and perimeter becomes 9 times greater. Its area becomes 6 times greater, and perimeter becomes 3 times greater. Its area becomes 3 times greater, and perimeter becomes 6 times greater.

OpenStudy (zepp):

Let the ratio be k Then: \(\begin{align} k&=3~(length)\\ k^2&=3^2=9~(Area) \\ k^3&=3^3=27~(Volume) \\ \end{align}\)

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

i understanddd ! :D

OpenStudy (zepp):

You are welcome! :)